OpenStudy (anonymous):

The number of kilometers Tony travels in a canoe, d, varies directly with the amount of time spent in the canoe, t. When Tony canoes for 2.25 h, he travels 9 km. Which equation shows this direct linear variation? a)d=2t b)d=2.25t c)d=4t D)d=4 + t

OpenStudy (anonymous):

do you know about direct and indirect variation?

OpenStudy (anonymous):

here the distance d varies directly with the time t so when d increases and t also increases

OpenStudy (anonymous):

all you need to do is to find their ratio of increase

OpenStudy (anonymous):

\[\frac{d}{t} = \frac{distance}{time}= \frac{9}{2.25}\]

OpenStudy (anonymous):

since 2.25 * 4 = 9 you'll get

OpenStudy (anonymous):

any doubts?

OpenStudy (anonymous):

\[\frac{d}{t} = 4\] which gives us d = 4t

OpenStudy (anonymous):

so c is the correct option.